# SpotFormer: Multi-Scale Spatio-Temporal Transformer for Facial Expression Spotting

### v2

@@ Abstract
Facial expression spotting, identifying periods where facial expressions occur in a video, is a significant yet challenging task in facial expression analysis. The issues of irrelevant facial movements and the challenge of detecting subtle motions in micro-expressions remain unresolved, hindering accurate expression spotting. In this paper, we propose an efficient framework for facial expression spotting. First, we propose a Sliding Window-based multi-temporal-resolution Optical flow (SW-MRO) feature, which calculates multi-temporal-resolution optical flow of the input image sequence within compact sliding windows. The window length is tailored to perceive complete micro-expressions and distinguish between general macro- and micro-expressions. SW-MRO can effectively reveal subtle motions while avoiding the optical flow being dominated by head movements. Second, we propose SpotFormer, a multi-scale spatio-temporal Transformer that simultaneously encodes spatio-temporal relationships of the SW-MRO features for accurate frame-level probability estimation. In SpotFormer, we use the proposed Facial Local Graph Pooling (FLGP) operation and convolutional layers to extract multi-scale spatio-temporal features. We show the validity of the architecture of SpotFormer by comparing it with several model variants. Third, we introduce supervised contrastive learning into SpotFormer to enhance the discriminability between different types of expressions. Extensive experiments on SAMM-LV, CAS(ME)2, and CAS(ME)3 show that our method outperforms state-of-the-art models, particularly in micro-expression spotting. Code will be available at https://github.com/KinopioIsAllIn/SpotFormer.
@@ 1 Introduction
Facial expressions can be broadly categorized into two groups: macro-expressions (MaEs) and micro-expressions (MEs). MaEs typically last from 0.5 to 4.0 seconds [nummenmaaekman] and are easily perceived by people due to their occurrence on a large facial area and high intensity [corneanu2016survey]. The analysis of MaEs is significant in various practical applications, such as sociable robots [fukuda2002facial], mental health [kopper1996experience], and virtual reality [facialvr]. In contrast, MEs generally last for less than 0.5 seconds [yan2013fast], and their perception is much more challenging due to their localized occurrence [bhushan2015study] and low intensity [wang2025micro]. Because of their involuntary nature, MEs are crucial in situations where people may attempt to conceal emotions or deceive others, such as in lie detection [ekman2009telling], medical care [endres2009micro], and national security [o2009police]. Therefore, both MaE and ME analysis play important roles in understanding human emotions and behaviors.
@@
Facial expression analysis generally consists of two steps: facial expression spotting (FES) and facial expression recognition. As the initial phase, FES aims to locate the onset and offset frames of MaE and ME intervals within video sequences, as illustrated in Fig. 1. The onset frame represents the beginning of an expression, while the offset frame indicates its end. However, FES is highly challenging due to several factors, including the difficulty in detecting subtle motions that exist in MEs, the presence of irrelevant motions (e.g., head movements and eye blinking) in long videos, and the difficulty of distinguishing between MaEs, MEs, and neutral expressions.
@@
In recent years, researchers have made significant progress in developing efficient algorithms for FES. Early studies employed traditional methods to extract hand-crafted features, such as optical flow in large sliding windows. Subsequently, they analyzed feature variations and detected expression intervals using threshold strategies [gan2020optical, he2020spotting, yuhong2021research, zhao2022rethinking]. More recently, with the development of deep learning, more and more researchers introduced learning-based methods to address the problem [verburg2019micro, wang2021mesnet, yap20223d, yin2023aware]. Additionally, due to the similarity between FES and temporal action localization (TAL), some researchers applied classical TAL frameworks to FES[yu2023lgsnet, GUO2023146] and introduced additional modules. These methods have achieved considerable improvements in MaE spotting performance.
@@
Despite significant progress made by previous studies, there are still some open questions that require exploration for further performance improvement, especially for ME spotting. The first question is how to extract more robust motion features. Traditional methods typically employ a large sliding window strategy to extract optical flow features and spot potential expression proposals within each sliding window [he2020spotting, guo2021magnitude]. However, the accuracy of using such a strategy is significantly impacted by head movement issues. In contrast, recent deep learning-based methods calculate optical flow between adjacent frames as motion features [yu2021lssnet, leng2022abpn]. Nevertheless, such an optical flow extraction method is insufficient to effectively reveal subtle motions present in MEs. Therefore, there is an urgent need to develop a strategy that can magnify motion information while mitigating the influence of head movement. Second, some methods [GUO2023146, yu2023lgsnet] applied general TAL frameworks to FES and refined them by introduing novel modules. However, these methods usually employ a pre-trained I3D[carreira2017quo] network to extract features from snippets based on the optical flow extracted between adjacent frames, which is unsuitable for revealing subtle motions as we mentioned before. Consequently, their performance, particularly in ME spotting, requires improvement. Third, some methods compute optical flow in specific regions of interest (ROIs) where facial expressions frequently happen to alleviate the influence of irrelevant motions. They then design a network to learn the extracted motion features [leng2022abpn, yin2023aware]. However, their proposed networks need a more comprehensive consideration of spatio-temporal relationships and multi-scale feature learning. This limitation restricts the representational capacity of the models.
@@
To address the issues mentioned above, we propose an efficient framework for facial expression spotting in untrimmed videos. First, we propose a sliding window-based multi-temporal-resolution optical flow (SW-MRO) feature to amplify motion information while mitigating head movement problems. Specifically, we calculate the optical flow between the first frame and each subsequent frame within compact sliding windows to generate multi-temporal-resolution optical flow features. Moreover, the temporal sliding window size is tailored to perceive complete subtle MEs and discern the differences between general MEs and MaEs. The SW-MRO feature enables our framework to achieve accurate frame-level apex or boundary (onset or offset) probability estimation and expression type classification. Second, we propose SpotFormer, a multi-scale spatio-temporal transformer to capture multi-scale spatial relationships and temporal variations among different facial parts across frames from the SW-MRO features. In SpotFormer, a facial local graph pooling (FLGP) operation is designed to extract multi-scale facial spatial features, while a learning-based temporal downsampling strategy is employed to extract multi-scale temporal features. These designs enhance the model’s understanding ability from low-level to high-level. The exploration of several model variants and extensive comparative experiments validate the effectiveness of SpotFormer. Third, we introduce supervised contrastive learning into our model to learn a finer discriminative feature representation. This contrastive learning approach alleviates the challenge of recognizing the classification boundaries between MaEs, MEs, and neutral expressions at the frame level.
@@
This paper is an extension of our prior conference publication [deng2024SpoTGCN]. The most crucial update is the proposal of SpotFormer. As other differences from the conference paper version, we make a more extensive overview and comparison of the related literature, particularly with regard to TAL. In addition, we present a more comprehensive qualitative and quantitative comparison to previous work in the experimental section. The contributions are listed as follows, where * represents new contributions of this paper:
@@
  • We propose SW-MRO, a sliding window-based multi-temporal-resolution optical flow feature for facial expression spotting, which can magnify the motion information while avoiding the optical flow being dominated by head movements. Extensive ablation studies demonstrate that such a carefully designed optical flow feature can assist the model not only in perceiving complete subtle MEs but also in discerning the differences between general MaEs and MEs.
@@ 2.1 Facial Expression Spotting
Early FES methods extracted appearance-based features, such as local binary patterns [moilanen2014spotting] and histogram of oriented gradients [davison2015micro]. These methods then employed machine learning algorithms and threshold strategies for expression spotting. Subsequently, See et al. [see2019megc] proposed the Facial Micro-Expressions Grand Challenge (MEGC), which introduced unified evaluation datasets and metrics for the FES task. The subsequent MEGC series [jingting2020megc2020, li2022megc2022, davison2023megc2023, see2024megc2024] have attracted increasing attention from researchers and significantly promoted the development of this field.
@@
In recent years, the mainstream approach for motion feature extraction shifted to optical flow. He et al. [he2020spotting] employed main directional maximal difference analysis [wang2017main] to detect facial movements and spot potential expression proposals. He [yuhong2021research] alleviated head movement problems by repeating face alignment using optical flow in the nose tip region. Guo et al. [guo2021magnitude] converted optical flow vectors into a polar coordinate representation and introduced a novel decision criterion accordingly. Recently, several researchers [zhao2022rethinking, yu2024micro] proposed novel methods to refine optical flow features or expression proposals, which further improved performance.
@@
With the development of deep learning, an increasing number of researchers have proposed various neural networks for feature learning. Zhang et al. [zhang2018smeconvnet] employed convolutional neural networks (CNNs) to extract features from video clips and spotted apex frames by analyzing the feature representations. Verburg et al. [verburg2019micro] extracted histogram of oriented optical flow features and employed an RNN to spot expression intervals. Liong et al. [liong2021shallow] introduced a three-stream CNN and employed pseudo-labeling techniques to facilitate the learning process. Yang et al. [yang2021facial] incorporated facial action unit information and concatenated various types of neural networks for feature learning. Leng et al. [leng2022abpn] calculated the main directional mean optical flow (MDMO) [liu2015main] optical flow features between adjacent frames and extended BSN [Lin_2018_ECCV] for frame-level probability estimation. Yin et al. [yin2023aware] extended ABPN [leng2022abpn] by introducing a graph convolutional network (GCN) to embed action unit (AU) label information into the extracted optical flow. Guo et al. [GUO2023146] proposed a multi-scale local transformer, similar to ActionFormer [zhang2022actionformer], and considered multi-scale temporal feature fusion for performance improvement. Yu et al. [yu2023lgsnet] applied A2Net [yang2020revisiting] to facial expression spotting and improved it by introducing attention modules. Zhang et al. [zhang2024multi] proposed a framework that integrates VideoMAE-based feature extraction, multi-scale candidate segment generation, and a multi-start-point optical flow filtering method for accurate ME and MaE detection.
@@
While current optical flow-based methods have significantly improved MaE spotting, their performance in ME spotting remains considerably lower. This is attributed to their extracted optical flow features, which cannot efficiently reveal subtle motions that exist in MEs. To solve this issue, we propose to extract a sliding window-based multi-temporal-resolution optical flow feature, which can effectively magnify these subtle motions and enhance the distinction between MaEs and MEs, thereby achieving accurate frame-level probability estimation and improving MaE and ME spotting performance.
@@ 2.2 Temporal Action Localization
Similar to FES, TAL aims to detect action intervals in general scene videos. Early works treated TAL as a 1D object detection task. Xu et al. [Xu_2017_ICCV] proposed a region convolutional 3D network to encode the video streams and locate the action proposals using an anchor-based method. Chao et al. [Chao_2018_CVPR] applied the faster R-CNN object detection framework to TAL by employing a multi-scale architecture to accommodate extreme variations in action duration. However, anchor-based methods require pre-defining a fixed number of anchors. The length and number of anchors need careful pre-definition, limiting flexibility and affecting the generalization ability of the proposed model.
In 2018, Lin et al. [Lin_2018_ECCV] proposed the Boundary-Sensitive Network (BSN), which was the first method to estimate frame-level probabilities and subsequently became the mainstream approach. Recently, Nag et al. [Nag_2023_CVPR] presented a novel Gaussian aware post-processing method for more accurate model inference, which models the start and end points of action instances with a Gaussian distribution to enable temporal boundary inference at a sub-snippet level.
Yang et al. [yang2020revisiting] were the first to directly regress the boundary offset for each temporal point, which solves the inflexibility and computation problems of anchor-based methods. Zhang et al. presented ActionFormer [zhang2022actionformer], a simple yet effective multi-scale transformer designed to further improve action localization performance.
@@
Firstly, motion information is usually obvious to detect in TAL videos, while in FES, the main challenge is to capture the subtle motion of MEs and suppress noises such as head movement and eye blinking. Since anchor-based methods and boundary regression-based methods typically require a large receptive field, as discussed before, extracting delicate optical flow becomes challenging. Therefore, in this paper, we design a frame-level probability-based framework following BSN [Lin_2018_ECCV].
Secondly, the scene usually changes in TAL videos, and the main challenge is to distinguish between the action scene and background change. In contrast, the scene during the whole video does not change much in FES, making it easier to capture vital motion information in certain facial areas that are strongly related to facial expressions. Therefore, in this paper, we represent the SW-MRO feature as spatio-temporal facial graphs and propose the SpotFormer to efficiently analyze the latent motion patterns.
Finally, TAL involves an action classification task for each proposal since their videos may involve hundreds of action classes. However, there are only two action classes (i.e., MaE and ME) in FES. Rather than focusing on action recognition, we aim to distinguish between MEs and MaEs. Therefore, in this paper, we emphasize perceiving MEs and analyzing the differences between MEs and MaEs within compact sliding windows, which inspired us to propose the SW-MRO feature and the supervised contrastive learning.
Figure 2: Overview of the proposed framework. (a) The data pre-processing module calculates the SW-MRO features; (b) the probability estimation module employs SpotFormer which takes optical flow features as input for frame-level apex or boundary probability estimation; (c) the post-processing module aggregates the probability maps from all frames and generates expression proposals.
@@ 2.3 Transformer
The Transformer was first proposed by Vaswani et al. [transformer] for natural language processing. In recent years, many researchers have explored the Transformer for computer vision and have shown great success in 2D image processing (e.g., ViT [dosovitskiy2021an], Swin Transformer [liu2021swin]) and 3D video understanding (e.g., ViViT [Arnab_2021_ICCV], Video Swin Transformer [Liu_2022_CVPR], ActionFormer [zhang2022actionformer]). Such success is attributed to their powerful self-attention mechanism and ability to model long-range relations. The self-attention mechanism allows the model to focus on critical features and ignore irrelevant information, enhancing performance and robustness. The long-range relation modeling enables efficient capture of context information across both spatial and temporal dimensions, contributing to a better understanding of the overall sequence. In this paper, we explore the potential of the graph-based Transformer for facial expression spotting and introduce SpotFormer, a novel model designed to enhance the accuracy of expression spotting.

### v3

@@ Abstract
Facial expression spotting, identifying periods where facial expressions occur in a video, is a significant yet challenging task in facial expression analysis. The issues of irrelevant facial movements and the challenge of detecting subtle motions in micro-expressions remain unresolved, hindering accurate expression spotting. In this paper, we propose an efficient framework for facial expression spotting. First, we propose a Compact Sliding-Window-based Multi-temporal-Resolution Optical flow (CSW-MRO) feature, which calculates multi-temporal-resolution optical flow of the input image sequence within compact sliding windows. The window length is tailored to perceive complete micro-expressions and distinguish between general macro- and micro-expressions. CSW-MRO can effectively reveal subtle motions while avoiding the optical flow being dominated by head movements. Second, we propose SpotFormer, a multi-scale spatio-temporal Transformer that simultaneously encodes spatio-temporal relationships of the CSW-MRO features for accurate frame-level probability estimation. In SpotFormer, we use the proposed Facial Local Graph Pooling (FLGP) operation and convolutional layers to extract multi-scale spatio-temporal features. We show the validity of the architecture of SpotFormer by comparing it with several model variants. Third, we introduce supervised contrastive learning into SpotFormer to enhance the discriminability between different types of expressions. Extensive experiments on SAMM-LV, CAS(ME)2, and CAS(ME)3 show that our method outperforms state-of-the-art models, particularly in micro-expression spotting. Code is available at https://github.com/KinopioIsAllIn/SpotFormer.
@@ 1 Introduction
Facial expressions can be broadly categorized into two groups: macro-expressions (MaEs) and micro-expressions (MEs). MaEs typically last from 0.5 to 4.0 seconds [13] and are easily perceived by people due to their occurrence on a large facial area and high intensity [8]. The analysis of MaEs is significant in various practical applications, such as sociable robots [18], mental health [26], and virtual reality [53]. In contrast, MEs generally last for less than 0.5 seconds [64] (mostly distributed around 306–328 ms [29]), and their perception is much more challenging due to their localized occurrence [2] and low intensity [58]. Because of their involuntary nature, MEs are crucial in situations where people may attempt to conceal emotions or deceive others, such as in lie detection [14], medical care [15], and national security [45]. Therefore, both MaE and ME analysis play important roles in understanding human emotions and behaviors.
@@
Facial expression analysis generally consists of two steps: facial expression spotting (FES) and facial expression recognition [79, 61]. As the initial phase, FES aims to locate the onset and offset frames of MaE and ME intervals within video sequences, as illustrated in Fig. 1. The onset frame represents the beginning of an expression, while the offset frame indicates its end. However, FES is highly challenging due to several factors, including the difficulty in detecting subtle motions that exist in MEs, the presence of irrelevant motions (e.g., head movements and eye blinking) in long videos, and the difficulty of distinguishing between MaEs, MEs, and neutral expressions.
@@
In recent years, researchers have made significant progress in developing efficient algorithms for FES. Early studies employed traditional methods to extract hand-crafted features, such as optical flow in large sliding windows. Subsequently, they analyzed feature variations and detected expression intervals using threshold strategies [19, 23, 77, 84]. More recently, with the development of deep learning, more and more researchers introduced learning-based methods to address the problem [55, 57, 70, 71]. Additionally, due to the similarity between FES and temporal action localization (TAL), some researchers applied classical TAL frameworks to FES[76, 20] and introduced additional modules. These methods have achieved considerable improvements in MaE spotting performance.
@@
Despite significant progress made by previous studies, there are still some open questions that require exploration for further performance improvement, especially for ME spotting. The first question is how to extract more robust motion features. Recent traditional methods typically employ a large sliding window strategy to extract optical flow features and spot potential expression proposals within each sliding window [23, 22]. However, the accuracy of using such a strategy is significantly impacted by head movement issues. In contrast, recent deep learning methods that employ optical flow for temporal modeling calculate optical flow between adjacent frames as motion features [75, 27]. Nevertheless, such an optical flow extraction method is insufficient to effectively reveal subtle motions present in MEs. Therefore, there is an urgent need to develop a strategy that can magnify motion information while mitigating the influence of head movement. Second, some methods [20, 76] applied general TAL frameworks to FES and refined them by introducing novel modules. However, these methods usually employ a pre-trained I3D[4] network to extract features from snippets based on the optical flow extracted between adjacent frames, which is unsuitable for revealing subtle motions as we mentioned before. Consequently, their performance, particularly in ME spotting, requires improvement. Third, some methods compute optical flow in specific regions of interest (ROIs) where facial expressions frequently happen to alleviate the influence of irrelevant motions. They then design a network to learn the extracted motion features [27, 71]. However, their proposed networks need a more comprehensive consideration of spatio-temporal relationships and multi-scale feature learning. This limitation restricts the representational capacity of the models.
@@
To address the issues mentioned above, we propose an efficient framework for facial expression spotting in untrimmed videos. First, we propose a compact sliding-window-based multi-temporal-resolution optical flow (CSW-MRO) feature to amplify motion information while mitigating head movement problems. Specifically, we calculate the optical flow between the first frame and each subsequent frame within compact sliding windows to generate multi-temporal-resolution optical flow features. Moreover, the temporal sliding window size is tailored to perceive complete subtle MEs and discern the differences between general MEs and MaEs. The CSW-MRO feature enables our framework to achieve accurate frame-level apex or boundary (onset or offset) probability estimation and expression type classification. Second, we propose SpotFormer, a multi-scale spatio-temporal transformer to capture multi-scale spatial relationships and temporal variations among different facial parts across frames from the CSW-MRO features. In SpotFormer, a facial local graph pooling (FLGP) operation is designed to extract multi-scale facial spatial features, while a learning-based temporal downsampling strategy is employed to extract multi-scale temporal features. These designs enhance the model’s understanding ability from low-level to high-level. The exploration of several model variants and extensive comparative experiments validate the effectiveness of SpotFormer. Third, we introduce supervised contrastive learning into our model to learn a finer discriminative feature representation. This contrastive learning approach alleviates the challenge of recognizing the classification boundaries between MaEs, MEs, and neutral expressions at the frame level.
@@
This paper is an extension of our prior conference publication [11]. The most crucial update is the proposal of SpotFormer. As other differences from the conference paper version, we make a more extensive overview and comparison of the related literature, particularly with regard to TAL. In addition, we present a more comprehensive qualitative and quantitative comparison to previous work in the experimental section. The contributions are listed as follows, where * represents new contributions of this paper:
@@
  • We propose CSW-MRO, a compact sliding-window-based multi-temporal-resolution optical flow feature for facial expression spotting, which can magnify the motion information while avoiding the optical flow being dominated by head movements. Extensive ablation studies demonstrate that such a carefully designed optical flow feature can assist the model not only in perceiving complete subtle MEs but also in discerning the differences between general MaEs and MEs.
@@ 2.1 Facial expression spotting
Early FES methods extracted appearance-based features, such as local binary patterns [43] and histogram of oriented gradients [10]. These methods then employed machine learning algorithms and threshold strategies for expression spotting. Subsequently, See et al. [51] proposed the Facial Micro-Expressions Grand Challenge (MEGC), which introduced unified evaluation datasets and metrics for the FES task. The subsequent MEGC series [24, 30, 9, 50] have attracted increasing attention from researchers and significantly promoted the development of this field.
@@
In recent years, the mainstream approach for motion feature extraction shifted to optical flow. He et al. [23] employed main directional maximal difference analysis [59] to detect facial movements and spot potential expression proposals. He [77] alleviated head movement problems by repeating face alignment using optical flow in the nose tip region. Guo et al. [22] converted optical flow vectors into a polar coordinate representation and introduced a novel decision criterion accordingly. Recently, several researchers [84, 74] proposed novel methods to refine optical flow features or expression proposals, which further improved performance.
@@
With the development of deep learning, an increasing number of researchers have proposed various neural networks for feature learning. Zhang et al. [83] employed convolutional neural networks (CNNs) to extract features from video clips and spotted apex frames by analyzing the feature representations. Verburg et al. [55] extracted histogram of oriented optical flow features and employed an RNN to spot expression intervals. Liong et al. [38] introduced a three-stream CNN and employed pseudo-labeling techniques to facilitate the learning process. Yang et al. [65] incorporated facial action unit information and concatenated various types of neural networks for feature learning. Leng et al. [27] calculated the main directional mean optical flow (MDMO) [39] optical flow features between adjacent frames and extended BSN [34] for frame-level probability estimation. Yin et al. [71] extended ABPN [27] by introducing a graph convolutional network (GCN) to embed action unit (AU) label information into the extracted optical flow. Guo et al. [20] proposed a multi-scale local transformer, similar to ActionFormer [78], and considered multi-scale temporal feature fusion for performance improvement. Yu et al. [76] applied A2Net [68] to facial expression spotting and improved it by introducing attention modules. Zhang et al. [82] proposed a framework that integrates VideoMAE-based feature extraction, multi-scale candidate segment generation, and a multi-start-point optical flow filtering method for accurate ME and MaE detection. Zou et al. [85] proposed a temporal state transition architecture based on state space models, which formulates spotting and recognition as a time-series regression problem rather than window-level classification.
@@
While current optical flow-based methods have significantly improved MaE spotting, their performance in ME spotting remains considerably lower. This is attributed to their extracted optical flow features, which cannot efficiently reveal subtle motions that exist in MEs. To solve this issue, we propose to extract a compact sliding-window-based multi-temporal-resolution optical flow feature, which can effectively magnify these subtle motions and enhance the distinction between MaEs and MEs, thereby achieving accurate frame-level probability estimation and improving MaE and ME spotting performance.
@@ 2.2 Temporal action localization
Similar to FES, TAL aims to detect action intervals in general scene videos. Early works treated TAL as a 1D object detection task. Xu et al. [62] proposed a region convolutional 3D network to encode the video streams and locate the action proposals using an anchor-based method. Chao et al. [5] applied the faster R-CNN object detection framework to TAL by employing a multi-scale architecture to accommodate extreme variations in action duration. However, anchor-based methods require pre-defining a fixed number of anchors. The length and number of anchors need careful pre-definition, limiting flexibility and affecting the generalization ability of the proposed model.
In 2018, Lin et al. [34] proposed the Boundary-Sensitive Network (BSN), which was the first method to estimate frame-level probabilities and subsequently became the mainstream approach. Recently, Nag et al. [44] presented a novel Gaussian aware post-processing method for more accurate model inference, which models the start and end points of action instances with a Gaussian distribution to enable temporal boundary inference at a sub-snippet level.
Yang et al. [68] were the first to directly regress the boundary offset for each temporal point, which solves the inflexibility and computation problems of anchor-based methods. Zhang et al. presented ActionFormer [78], a simple yet effective multi-scale transformer designed to further improve action localization performance.
@@
Firstly, motion information is usually obvious to detect in TAL videos, while in FES, the main challenge is to capture the subtle motion of MEs and suppress noises such as head movement and eye blinking. Since anchor-based methods and boundary regression-based methods typically require a large receptive field, as discussed before, extracting delicate optical flow becomes challenging. Therefore, in this paper, we design a frame-level probability-based framework following BSN [34].
Secondly, the scene usually changes in TAL videos, and the main challenge is to distinguish between the action scene and background change. In contrast, the scene during the whole video does not change much in FES, making it easier to capture vital motion information in certain facial areas that are strongly related to facial expressions. Therefore, in this paper, we represent the CSW-MRO feature as spatio-temporal facial graphs and propose the SpotFormer to efficiently analyze the latent motion patterns.
Finally, TAL involves an action classification task for each proposal since their videos may involve hundreds of action classes. However, there are only two action classes (i.e., MaE and ME) in FES. Rather than focusing on action recognition, we aim to distinguish between MEs and MaEs. Therefore, in this paper, we emphasize perceiving MEs and analyzing the differences between MEs and MaEs within compact sliding windows, which inspired us to propose the CSW-MRO feature and the supervised contrastive learning.
Figure 2: Overview of the proposed framework. (a) The data pre-processing module calculates the CSW-MRO features; (b) the probability estimation module employs SpotFormer which takes optical flow features as input for frame-level apex or boundary probability estimation; (c) the post-processing module detects pseudo-apex frames, generates expression proposals by searching for onset and offset frames around each pseudo-apex, and filters redundant proposals using NMS.
@@ 2.3 Transformer
The Transformer was first proposed by Vaswani et al. [54] for natural language processing. In recent years, many researchers have explored the Transformer for computer vision and have shown great success in 2D image processing (e.g., ViT [12], Swin Transformer [40]) and 3D video understanding (e.g., ViViT [1], Video Swin Transformer [41], ActionFormer [78]). In recent years, several efforts have been devoted to investigating Transformers for macro- or micro-expression recognition [33, 60].
Such success is attributed to their powerful self-attention mechanism and ability to model long-range relations. The self-attention mechanism allows the model to focus on critical features and ignore irrelevant information, enhancing performance and robustness. The long-range relation modeling enables efficient capture of context information across both spatial and temporal dimensions, contributing to a better understanding of the overall sequence. In this paper, we explore the potential of the graph-based Transformer for facial expression spotting and introduce SpotFormer, a novel model designed to enhance the accuracy of expression spotting.